answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

HOW you can simulat the poisson random measure?

A simple way of generating poisson random numbers is given by Knuth, in his book "The Art of Computer Programming" in the chapter on Seminumerical Algorithms. I can't quote it directly here, due to copyright and plagarism limitations, but the Related Link below contains a reference to it.

What is the result of the combination when n is zero?

you r mad we will n0t answer it and just shut up . this site is for answers not for answering. fools.

How many equal groups of cards can be made from a standard deck of 52 cards?

52/1 = 52

52/2 = 26

52/4 = 13

52/13 = 4

52/26 = 2

52/52 = 1

6 groups. 1 group of 52 counts, right?

Why are the kings called cowboys in a deck of cards?

The only reason known is simply the reputation that early time cowboys had for playing Poker. There may also be a phonetic connection between the sound of the 'k' in king and the 'c' in cowboy.

What is the advantage of a one-tailed test?

A one tailed test allows you to test a one-sided hypothesis.

How many possible outcomes are there rolling six sided cube twice?

  • If each throw is unique, then there are 36 distinct outcomes(6 x 6 = 36).
  • If each throw is not unique (you count 1-4 same as 4-1) then 21 different possible.
  • If you're just interested in the sums, then 11 different possible (1+1 =2 up to 6+6=12)

But any time you want to calculate probabilities with two die throws, you need to use the 36 distinct outcomes.

How do you convert a z-score into an x-score?

Suppose m is the mean and s the standard deviation of a random variable, X, which is normally distributed. Then, given Z,

X = m + sZ

Which car will move first if their are 15 marbles while the other has 10 marbles?

I presume you mean they're on a slope. In which case, the car with the most marbles will move first, because it's heavier.

What is the differences between equal probability and equal width in measurement?

One is a measure of probability, the other is a measure of width! And neither is the same as equal age, or equal loudness!

One is a measure of probability, the other is a measure of width! And neither is the same as equal age, or equal loudness!

One is a measure of probability, the other is a measure of width! And neither is the same as equal age, or equal loudness!

One is a measure of probability, the other is a measure of width! And neither is the same as equal age, or equal loudness!

In an experiment, a biased coin is thrown ten times. If the probability of obtaining a head on any throw is 0.4 find the probability of obtaining?

Mr. Palmer moved to Canada in 1974 where he continued to teach and write. He died there on June 16, 2013. Related Documents.

Right as a fraction and probability of rolling a three on a six sided dice?

The probability of NOT rolling a 3 with one die is 5/6 so the probability of NOT rolling a 3 with a roll of two dice is 25/36. The probability of rolling at least one 3 is 1–25/36=11/36, a bit less than 1/3.

How many sides are there in a die?

A standard six-sided die has six sides, as the name suggests.

However, dice can have four or more sides; the term die(or dice) does not imply a particular number of sides.

Some board games include dice with unusual shape and a number of sides different from six. Mathematicians also often conduct thought-experiments based on a notion of n-sided dice.

The probability of this kind of event is 0?

If the probability is 0, the event has no chance of happening.